Kieso Company borrowed $780,000 for six months. The annual interest rate on the loan was 11%. Kieso's fiscal year ends on December 31. Kieso borrowed the $780,000 one month prior to the start of its current fiscal year and paid back the $780,000 plus interest five months into its current fiscal year.

How much interest expense, if any, would Kieso report at the end of its last fiscal year and at the end of its current fiscal year?

Last year:

Current Year:
Answer #1

Last year = 780,000 * 11%* 1/12 = 7150 [ for one month in the previous fiscal year]

Current year = 780,000 * 11%* 5/12 = 35750 [ for five months in the current fiscal year]